### v1

@@ 4.2 A High-Frequency Narrative Approach
While the approach employed in the previous section has several advantages, including its simplicity, there are a some drawbacks as well. Election betting markets are noisy and contain a lot of price movements that may not be well explained by actual shifts in election probabilities. Additionally, even if these markets perfectly capture reality, it would be optimistic to believe that investors and other economic agents react to subtle changes in the election outlook that happen on a day-to-day basis. Contrarily, investors are likely to react to newsworthy events that lead to large changes in the election outlook. With this in mind, here I construct an alternative measure that considers only changes in election probabilities around notable narrative events, and repeat the analysis from the previous section.
Table C2 provides a list of dates and associated events from the 2000-2024 elections. Each presidential and vice presidential debate is included along with other exogenous events that occur in conjunction with notable shifts in election probabilities. [^21] Each Democratic and Republican national conventions are also included, since these events are widely televised and scrutinized by political analysts. In practice, constructing a narrative-based shock series requires making a choice about how tight of a window around each date to attribute changes in election probabilities. It is standard in the high-frequency identification literature to use one-day or even intraday windows. For this application, it is not clear that the tightest possible window (of one day) is advantageous. As discussed in Section 3.3, in some cases, election probabilities react somewhat sluggishly to events, which is captured by the fact that many important dates are followed by a cluster of shocks of the same sign rather than one discrete movement. However, larger windows increase the likelihood of contaminating the measure with noise. With this in mind, for my baseline narrative shock, I consider a 5-day window [^22] In unreported results I also consider 3-day and 1-day windows with no qualitative affect on the results. and let:
@@
where $\mathbbm{1}_{N}(t)$ is an indicator function equal to one for each the dates given in Table C2. Note that the use of the original shock measure rather than the raw election probabilities reduces the potential for confounding the narrative shocks with financial and economic ones. Figure C2 plots $\text{Shock}_{t}^{\text{N}}$ both for the full sample and separately for each election window. Similarly to $\text{Shock}_{t}$, the largest shocks tend to be on election day (particularly in 2016 and 2024), though for the majority of years the absolute sum of non-election day values is larger than the election day value. As in the baseline series, shocks in favor of Republicans and Democrats tend to align with the types of events one would expect given the narrative information around each selected date. See Appendix C for more discussion.
With the alternative series in hand, I re-estimate equation (4), replacing $\text{Shock}_{t}$ with $\text{Shock}_{t}^{\text{N}}$.
@@
Figure 7 plots the impulse responses of each sectoral asset price to a 10pp narrative “Republican shock.” The results largely mimic that of Figure 6. The energy response tends to be a bit smaller in the medium-term, while the clean energy and defense responses are larger, albeit with wider confidence intervals at longer horizons. Despite these differences, the estimated responses are of the same order of magnitude as in the baseline results, reinforcing the takeaways from the previous section.
@@ 4.3 Labor Markets
where $y^{i}_{t}$ is the log private employment of industry $i$, $\Delta y^{i}_{t+h}$ is an $h$-month ahead long difference, $\overline{\text{Shock}}_{t}$ is the monthly sum of $\text{Shock}_{t}$, and $\mathbf{W}^{i}_{t}$ is a vector containing $y^{i}_{t}$, the unemployment rate, and 1-month differences in log CPI, log PCE, and log industrial production. $\theta^{i}_{h}$ are the coefficients of interest. Aside from aggregate employment, the set of industries I consider are oil drilling and extraction, mining and quarrying, clean energy generation, aerospace manufacturing, and ship and tank manufacturing. For each horizon and industry, I estimate equation (6) on the data from January 2002 through March 2024, omitting observations where $t+h$ is between March and December 2020 to avoid complications related to the COVID-19 period. [^23] Lenza and Primiceri (2022) argue that this is an acceptable strategy to avoid the distortion of parameter estimates in VAR models. For more details about the set of industries see Table D1.
Figure 8 plots the estimated 0-12 month-ahead industrial employment response to a 10pp election shock along with the 68 and 90 percent confidence bands. [^24] Standard errors are estimated in the same fashion as in the previous section. The first panel gives the total employment response. It is not ex ante obvious whether total employment should favor one of the two political parties. For instance, Republicans generally favor lower taxes while Democrats favor larger amounts of federal spending, both of which could stimulate the economy. Nonetheless, the employment response to an election shock remains close to zero and statistically insignificant throughout.
Industries associated with the energy sector respond positively to a Republican shock. Oil drilling and extraction employment increases with a slight lag with point estimates peaking at roughly 1.6 percent 10 months after the shock. Similarly, mining and quarrying employment increases contemporaneously, with the response peaking 7 months after the shock at roughly 1 percent. On the other hand, jobs in clean energy generation display little response in the short-term but decrease by approximately 2 percent 12 months after the shock. Overall, the energy and clean energy sectors respond in the expected direction given the discussion in Section 2, albeit with varying lags and magnitudes. As in the stock markets application I interpret 0-3 month effects as being driven purely by the election news channel, while responses after the 3-month mark may be driven by a combination of news and actual policy implementation.
@@
Aerospace manufacturing, which makes up approximately 76% of defense-related manufacturing jobs in the sample does not respond. On the other hand, ship and tank manufacturing shows a small contemporaneous increase, albeit without lasting effects. [^25] Figure D5 reveals that the effects on tank manufacturing, which makes up a relatively smaller share of jobs, are more persistent. Taken together, there is some evidence that election shocks have an impact on defense manufacturing employment. Given the significant lobbying of defense contractors to both political parties, it is unsurprising that these effects are muted relative to jobs in energy and clean energy.
@@ 5 Conclusion
Studying the affects of government policies on the macroeconomy generally requires one to measure the response of relevant economic variables to the policies’ implementation. However, previous work has argued that much of the response often takes place in anticipation of policy implementation, complicating researchers’ task. This paper provides evidence of these anticipation effects in the context of U.S. presidential elections. Election outcomes and changes in election probabilities throughout the election period have strong affects on stock markets. Relative to the Democratic party, changes in favor of the Republican party are correlated with future policies that expand the energy and defense sectors and shrink the clean energy sector. In response to these changes, energy and defense asset prices increase while clean energy prices decrease. Similarly, employment in industries associated with each of these sectors change in the short- to medium-term. In particular, oil, mining, and some types of defense manufacturing jobs increase in response to Republican shocks while clean energy generation jobs decrease. One potential avenue for future research is to study whether the labor market response to these shocks is driven by the stock market response, firm expectations, or another mechanism. Future work may also contribute by disentangling the cause of medium- and long-term responses, which may be confounded with policy implementation.
@@ Appendix A Political Polarization in Energy Policy
This section provides additional information on energy policy implementation of Republican and Democratic administrations. As discussed in Section 2, over the past several decades Republican administrations have generally passed legislation that stimulates the expansion of oil, gas, and coal and the contraction of clean energy sources relative to Democratic administrations. One of the most direct ways that the Federal government benefits (or inhibits) the traditional energy sector is by its approval of pipeline construction. Pipelines are used to transport natural gas from production areas to storage facilities and consumers. Figure A1 plots the miles of pipe approved by the Federal government in each year from 2000 through 2024. Red vertical bars indicate the start of a new Republican administration, while Blue bars indicate the start of a new Democratic administration. Overall, Republicans approved 180% more miles of pipe on average per year throughout the sample. Notably, the start of each Republican administration is associated with a large spike in approval. In practice, this may be due to the approval of projects that were inhibited by their Democratic predecessors. For instance, within Trump’s first month in office in 2017 he approved the Dakota pipeline, which is a 1,172 mile long pipeline for which construction was blocked in 2016 by the Obama administration. Overall, this Figure provides evidence that Democrats have been more likely to inhibit the transportation of gas than Republicans in recent decades.
@@
Aside from pipeline construction, there are a host of actions taken by recent administrations that approach the energy and clean energy sectors from different angles. Perhaps most symbolically, recent administrations have estimated the social cost of carbon, which is a measure intended to help policymakers weigh the economic benefits and drawbacks of energy related policies that might increase or decrease emissions. This measure was originally introduced under the Obama administration at \$43 a per metric ton of C02. The Trump administration’s estimate was \$3 - \$5 a ton, 88% - 93%(!) lower than the original estimate. Finally, the Biden administration’s estimate was approximately \$51 a ton. These vast differences indicate that, at least for the past few administrations, there is a stark divide between the way Republicans and Democrats view the necessity of reducing emissions. These differences have also led to a notable divergence in the types of measures taken by each party. Table A1 provides a comprehensive list of significant energy-related actions taken by recent administrations. Actions include, but are not limited to, international emissions reduction agreements, direct investments, tax incentives, vehicle emissions standards, loans, and royalties and leases for drilling on public lands.
Table 1: *
@@ Appendix B Iowa Electronic Markets Data
This section provides supplementary information on the data taken from Iowa Electronic Markets (IEM). IEM is operated by faculty at the University of Iowa Henry B. Tippie College of Business for research purposes. Throughout, I use prices from their winner take all U.S. presidential election markets. For each election traders may buy assets titled DEMYY_WTA or REPYY_WTA, where YY represents the year of the election. For example, a DEM24_WTA contract pays \$1 if the Democratic party nominee receives the majority of the popular vote cast in 2024, and \$0 otherwise. Thus, through the “wisdom of the masses,” the price of each contract on any given day is a forecast of the probability that an event will happen. To construct the probabilities that each party nominee wins the election I take the last price for the two party contracts from each day in the sample, and normalize the sum of the two prices to one. [^26] In practice the sum of the two prices are usually within a few hundredths of one, though the existence of 3rd party contracts and low trade volumes on certain days sometimes leads to slight deviations. For instance, on November 1, 2000, the last prices of the DEM00_WTA and REP00_WTA contracts were \$0.348 and \$0.668, respectively. After normalization the implied probabilities for the Democratic and Republican candidates were 34.3% and 65.7%, respectively. Data start May 1st, 2000 for the 2000 election, June 1st, 2004 for the 2004 election, and by at least January 1st of the election year for the remaining elections in the sample. [^27] Throughout I only consider data starting in January of the election year. Figure B1 plots the IEM implied probabilities for each election period in the sample. Vertical bars in each subplot indicate the day of the election.
@@ Appendix C Election Shocks and Events
As discussed in Section 4.2, another approach to constructing the election shock series is to use only changes near prominent election events. Table C2 lists the events considered in this paper, and provides a description of each event where relevant. For each election I include each presidential and vice presidential debate, as well as any scandals that are associated with changes in election probabilities through the altering of public opinion on one of the candidates. Many of these events were already discussed in Section 3.3. Some other recent notable events are the release of the Access Hollywood tape in the 2016 election, and the White House lockdown on May 29, 2020 amidst protests in response to George Floyd’s death in an encounter with a police officer.
Table 2: *
@@ D.1 Stock Market Effects
weighting by trade volumes as in Section 4.1. Shocks are normalized to increase $\pi^{R}_{t}$ by 10pp. Overall, the initial effects of the shock are persistent with the estimated coefficients at the furthest horizon still remaining statistically greater than zero and point estimates only dropping by roughly 2pp. These dynamics are partly mechanical: $\pi^{R}_{t}$ persistently goes to zero or one in the period just after an election. Thus, a shock at time $t$ will trivially be associated with persistent effects at horizon $h$ when $t+h$ takes places after election day. In unreported results I re-estimate the dynamic response of $\pi^{R}_{t}$ without including horizons that take place after the election and the response remains persistent, albeit with more volatile point estimates and larger confidence bands.
@@
where $\text{Elec}_{t}$ equals -1 (1) on dates that a Democrat (Republican) wins a presidential election and 0. As in $\text{Shock}_{t}$, this variable treats Democratic and Republican victories symmetrically, but ignores the probabilities assigned to each event. In the main text I refer to this methodology as the “crude” approach. The advantage of this approach is its simplicity. However there are some substantial drawbacks. First, the approach misses any innovations in election probabilities leading up to election day, reducing the sample of non-zero innovations to seven. Second, it weights each election day outcome the same regardless of the probability of the event. For instance, Obama’s expected victory in 2008 is treated symmetrically to Trump’s unexpected victory in 2016. If election probabilities are already factored into the price of sectoral stock prices, as narrative evidence provided in Figure 3 suggests, this methodology fails to account for important anticipatory affects. Figure D2 plots the response of each sectoral stock price to a Republican election victory. Each price responds in the expected direction contemporaneously, however in the medium-term the point estimates for Energy and Defense are near zero, or even below zero, albeit to a nonsignificant degree. The response of clean energy has a comparable shape to that of the main results displayed in Figure 6. Overall, the failure to find persistent effects in Energy and Defense indicates that it is important to consider the information provided by election probabilities throughout an election cycle.
@@
As a robustness check to the main results 4.1, I also take a one-step approach to estimating the effects of an exogenous increase $\pi^{R}_{t}$ without estimating $\text{Shock}^{R}_{t}$. To do so I estimate the equation:
@@
where each variable is defined the same way as Sections 3 and 4, and the coefficient of interest is $\alpha^{i}_{1,0}$, i.e., the contemporaneous effect of a change in $\pi^{R}_{t}$. The identifying assumption is that after controlling for the other included regressors, changes to $\pi^{R}_{t}$ are exogenous to $\Delta y^{i}_{t+h}$. I estimate equation D3 using weighted OLS. Figure D2 plots the response of each sectoral stock price to a 10pp increase to $\pi^{R}_{t}$. The responses largely mimic those of the main results plotted in Figure 6.

### v2

@@ 4.2 A High-Frequency Narrative Approach
While the approach employed in the previous section has several advantages, including its simplicity, there are a some drawbacks as well. Election betting markets are noisy and contain a lot of price movements that may not be well explained by actual shifts in election probabilities. Additionally, even if these markets perfectly capture reality, investors and other economic agents may not react to subtle changes in the election outlook that happen on a day-to-day basis. Contrarily, investors are more likely to react to newsworthy events that lead to large changes in the election outlook. With this in mind, here I construct an alternative measure that considers only changes in election probabilities around notable narrative events, and repeat the analysis from the previous section.
Table LABEL:tabc2 provides a list of dates and associated events from the 2000-2024 elections. Each presidential and vice presidential debate is included along with other exogenous events that occur in conjunction with notable shifts in election probabilities. [^20] Each Democratic and Republican national conventions are also included, since these events are widely televised and scrutinized by political analysts. In practice, constructing a narrative-based shock series requires making a choice about how tight of a window around each date to attribute changes in election probabilities. It is standard in the high-frequency identification literature to use one-day or even intraday windows. For this application, it is not clear that the tightest possible window (of one day) is advantageous. As discussed in Section 3.3, in some cases, election probabilities react somewhat sluggishly to events, which is captured by the fact that many important dates are followed by a cluster of shocks of the same sign rather than one discrete movement. However, larger windows increase the likelihood of contaminating the measure with noise. With this in mind, for my baseline narrative shock, I consider a 5-day window [^21] In unreported results I also consider 3-day and 1-day windows with no qualitative affect on the results. and let:
@@
where $\mathbbm{1}_{N}(t)$ is an indicator function equal to one for each the dates given in Table LABEL:tabc2. Note that the use of the original shock measure rather than the raw election probabilities reduces the potential for confounding the narrative shocks with financial and economic ones. Figure C2 plots $\text{Shock}_{t}^{\text{N}}$ both for the full sample and separately for each election window. Similarly to $\text{Shock}_{t}$, the largest shocks tend to be on election day (particularly in 2016 and 2024), though for the majority of years the absolute sum of non-election day values is larger than the election day value. As in the baseline series, shocks in favor of Republicans and Democrats tend to align with the types of events one would expect given the narrative information around each selected date. See Appendix C for more discussion.
With the alternative series in hand, I re-estimate equation (4), replacing $\text{Shock}_{t}$ with $\text{Shock}_{t}^{\text{N}}$. Figure 7 plots the impulse responses of each sectoral asset price to a 10pp narrative “Republican shock.” The results largely mimic that of Figure 6. The energy response tends to be a bit smaller in the medium-term, while the clean energy and defense responses are larger, albeit with wider confidence intervals at longer horizons. Despite these differences, the estimated responses are of the same order of magnitude as in the baseline results, reinforcing the takeaways from the previous section.
@@ 4.3 Labor Markets
where $y^{i}_{t}$ is the log private employment of industry $i$, $\Delta y^{i}_{t+h}$ is an $h$-month ahead long difference, $\overline{\text{Shock}}_{t}$ is the monthly sum of $\text{Shock}_{t}$, and $\mathbf{W}^{i}_{t}$ is a vector containing $y^{i}_{t}$, the unemployment rate, and 1-month differences in log CPI, log PCE, and log industrial production. $\theta^{i}_{h}$ are the coefficients of interest. Aside from aggregate employment, the set of industries I consider are oil drilling and extraction, mining and quarrying, clean energy generation, aerospace manufacturing, and ship and tank manufacturing. Table D1 provides details about the set of industries. For each horizon and industry, I estimate equation (6) on the data from January 2002 through March 2024.222222I omit observations where $t+h$ is between March and December 2020 to avoid complications related to the COVID-19 period. See Lenza and Primiceri (2022) for a detailed discussion of this issue.
Figure 8 plots the estimated 0-12 month-ahead industrial employment response to a 10pp election shock along with the 68 and 90 percent confidence bands. [^23] Standard errors are estimated in the same fashion as in the previous section. The first panel gives the total employment response, which remains close to zero and statistically insignificant throughout.
@@
Industries associated with the energy sector respond positively to a Republican shock. Oil drilling and extraction employment increases with a slight lag and point estimates peaking at roughly 1.6 percent 10 months after the shock. Similarly, mining and quarrying employment increases contemporaneously, with the response peaking 7 months after the shock at roughly 1 percent. On the other hand, jobs in clean energy generation display little response in the short-term but decrease by approximately 2 percent 12 months after the shock. Overall, the energy and clean energy sectors respond in the expected direction given the discussion in Section 2, albeit with varying lags and magnitudes. As in the stock markets application I interpret 0-3 month effects as being driven purely by the election news channel, while responses after the 3-month mark may be driven by a combination of news and actual policy implementation.
Aerospace manufacturing, which makes up approximately 76 percent of defense-related manufacturing jobs in the sample does not respond. On the other hand, ship and tank manufacturing shows a small contemporaneous increase, albeit without lasting effects. [^24] Figure D5 reveals that the effects on tank manufacturing, which makes up a relatively smaller share of jobs, are more persistent. Taken together, there is some evidence that election shocks have an impact on defense manufacturing employment. Given the significant lobbying of defense contractors to both political parties, it is unsurprising that these effects are muted relative to jobs in energy and clean energy.
@@ 5 Conclusion
Studying the effects of government policies on the macroeconomy generally requires one to measure the response of relevant economic variables to the policies’ implementation. However, previous work has argued that much of the response often takes place in anticipation of policy implementation, complicating researchers’ task. This paper provides evidence of these anticipation effects in the context of U.S. presidential elections. Election outcomes and changes in election probabilities throughout the election period have strong effects on stock markets. Relative to the Democratic party, changes in favor of the Republican party are correlated with future policies that expand the energy and defense sectors and shrink the clean energy sector. In response to these changes, energy and defense asset prices increase while clean energy prices decrease. Similarly, employment in industries associated with each of these sectors change in the short- to medium-term. In particular, oil, mining, and some types of defense manufacturing jobs increase in response to Republican shocks while clean energy generation jobs decrease. One potential avenue for future research is to study whether the labor market response to these shocks is driven by the stock market response, firm expectations, or another mechanism. Future work may also contribute by disentangling news-driven effects from policy implementation in the medium- and long-term.
@@ Appendix A Political Polarization in Energy Policy
This section provides additional information on energy policy implementation of Republican and Democratic administrations. As discussed in Section 2, over the past several decades Republican administrations have generally passed legislation that stimulates the expansion of oil, gas, and coal and the contraction of clean energy sources relative to Democratic administrations. One of the most direct ways that the Federal government benefits (or inhibits) the traditional energy sector is by its approval of pipeline construction. Pipelines are used to transport natural gas from production areas to storage facilities and consumers. Figure A1 plots the miles of pipe approved by the Federal government in each year from 2000 through 2024. Red vertical bars indicate the start of a new Republican administration, while Blue bars indicate the start of a new Democratic administration. Overall, Republicans approved 180 percent more miles of pipe on average per year throughout the sample. Notably, the start of each Republican administration is associated with a large spike in approval. In practice, this may be due to the approval of projects that were inhibited by their Democratic predecessors. For instance, within Trump’s first month in office in 2017 he approved the Dakota pipeline, which is a 1,172 mile long pipeline for which construction was blocked in 2016 by the Obama administration. Overall, this figure provides evidence that Democrats have been more likely to inhibit the transportation of gas than Republicans in recent decades.
@@
Aside from pipeline construction, there are a host of actions taken by recent administrations that approach the energy and clean energy sectors from different angles. Perhaps most symbolically, recent administrations have estimated the social cost of carbon, which is a measure intended to help policymakers weigh the economic benefits and drawbacks of energy related policies that might increase or decrease emissions. This measure was originally introduced under the Obama administration at \$43 a per metric ton of C02. The Trump administration’s estimate was \$3 - \$5 a ton, 88 - 93 percent(!) lower than the original estimate. Finally, the Biden administration’s estimate was approximately \$51 a ton. These vast differences indicate that, at least for the past few administrations, there is a stark divide between the way Republicans and Democrats view the necessity of reducing emissions. These differences have also led to a notable divergence in the types of measures taken by each party. Table LABEL:taba1 provides a comprehensive list of significant energy-related actions taken by recent administrations. Actions include, but are not limited to, international emissions reduction agreements, direct investments, tax incentives, vehicle emissions standards, loans, and royalties and leases for drilling on public lands.
@@ Appendix B Iowa Electronic Markets Data
This section provides supplementary information on the data taken from Iowa Electronic Markets (IEM). IEM is operated by faculty at the University of Iowa Henry B. Tippie College of Business for research purposes. Throughout, I use prices from their winner take all U.S. presidential election markets. For each election traders may buy assets titled DEMYY_WTA or REPYY_WTA, where YY represents the year of the election. For example, a DEM24_WTA contract pays \$1 if the Democratic party nominee receives the majority of the popular vote cast in 2024, and \$0 otherwise. Thus, through the “wisdom of the masses,” the price of each contract on any given day is a forecast of the probability that an event will happen. To construct the probabilities that each party nominee wins the election I take the last price for the two party contracts from each day in the sample, and normalize the sum of the two prices to one. [^25] In practice the sum of the two prices are usually within a few hundredths of one, though the existence of 3rd party contracts and low trade volumes on certain days sometimes leads to slight deviations. For instance, on November 1, 2000, the last prices of the DEM00_WTA and REP00_WTA contracts were \$0.348 and \$0.668, respectively. After normalization the implied probabilities for the Democratic and Republican candidates were 34.3 and 65.7 percent, respectively. Data start May 1st, 2000 for the 2000 election, June 1st, 2004 for the 2004 election, and by at least January 1st of the election year for the remaining elections in the sample. [^26] Throughout I only consider data starting in January of the election year. Figure B1 plots the IEM implied probabilities for each election period in the sample. Vertical bars in each subplot indicate the day of the election.
@@ Appendix C Election Shocks and Events
As discussed in Section 4.2, another approach to constructing the election shock series is to use only changes near prominent election events. Table LABEL:tabc2 lists the events considered in this paper, and provides a description of each event where relevant. For each election I include each presidential and vice presidential debate, as well as any scandals that are associated with changes in election probabilities through the altering of public opinion on one of the candidates. Many of these events were already discussed in Section 3.3. Some other recent notable events are the release of the Access Hollywood tape in the 2016 election, and the White House lockdown on May 29, 2020 amidst protests in response to George Floyd’s death in an encounter with a police officer.
@@ D.1 Stock Market Effects
weighting by trade volumes as in Section 4.1. Shocks are normalized to increase $\pi^{R}_{t}$ by 10pp. Overall, the initial effects of the shock are persistent with the estimated coefficients at the furthest horizon still remaining statistically greater than zero and point estimates only dropping by roughly percent . These dynamics are partly mechanical: $\pi^{R}_{t}$ persistently goes to zero or one in the period just after an election. Thus, a shock at time $t$ will trivially be associated with persistent effects at horizon $h$ when $t+h$ takes places after election day. In unreported results I re-estimate the dynamic response of $\pi^{R}_{t}$ without including horizons that take place after the election and the response remains persistent, albeit with more volatile point estimates and larger confidence bands.
@@
where $\text{Elec}_{t}$ equals -1 (1) on dates that a Democrat (Republican) wins a presidential election and 0. As in $\text{Shock}_{t}$, this variable treats Democratic and Republican victories symmetrically, but ignores the probabilities assigned to each event. In the main text I refer to this methodology as the “crude” approach. The advantage of this approach is its simplicity. However there are some substantial drawbacks. First, the approach misses any innovations in election probabilities leading up to election day, reducing the sample of non-zero innovations to seven. Second, it weights each election day outcome the same regardless of the probability of the event. For instance, Obama’s expected victory in 2008 is treated symmetrically to Trump’s unexpected victory in 2016. If election probabilities are already factored into the price of sectoral stock prices, as narrative evidence provided in Figure 3 suggests, this methodology fails to account for important anticipatory effects. Figure D2 plots the response of each sectoral stock price to a Republican election victory. Each price responds in the expected direction contemporaneously, however in the medium-term the point estimates for Energy and Defense are near zero, or even below zero, albeit to a nonsignificant degree. The response of clean energy has a comparable shape to that of the main results displayed in Figure 6. Overall, the failure to find persistent effects in Energy and Defense indicates that it is important to consider the information provided by election probabilities throughout an election cycle.
@@
As a robustness check to the main results 4.1, I also take a one-step approach to estimating the effects of an exogenous increase $\pi^{R}_{t}$ without estimating $\text{Shock}_{t}$. To do so I estimate the equation:
@@
where each variable is defined the same way as Sections 3 and 4, and the coefficient of interest is $\alpha^{i}_{1,0}$, i.e., the contemporaneous effect of a change in $\pi^{R}_{t}$. As opposed to the baseline specification, this approach does not suffer from a potential generated regressor problem. The identifying assumption is that after controlling for the other included regressors, changes to $\pi^{R}_{t}$ are exogenous to $\Delta y^{i}_{t+h}$. I estimate equation (D3) using weighted OLS. Figure D3 plots the response of each sectoral stock price to a 10pp increase to $\pi^{R}_{t}$. The responses largely mimic those of the main results plotted in Figure 6.